# Decentralization of Ethereum’s Builder Market

### v1

@@ 2.1. MEV-Boost Auctions
MEV.   Miner/Maximal Extractable Value (MEV) (Daian et al., 2020) refers to the profit that privileged players (e.g., validators) can earn by including, excluding, and reordering the transactions. MEV poses a threat to decentralization, as explained in the introduction. There are two schools of thought on mitigating the negative impacts of MEV. The first one is referred to as MEV democratization (Flashbots, 2022b), the main idea of which is to facilitate MEV extraction and level the playground. MEV-Boost (Flashbots, 2024), to be introduced below, is the primary example. Another approach is to enforce certain ordering policies (e.g., FIFO) to prevent order manipulation, such as (Kelkar et al., 2020, 2023). We focus on MEV-Boost because it is widely used.
@@
PoS Ethereum.   Before introducing background on MEV-Boost auctions, we first cover the relevant concepts in Ethereum’s Proof of Stake-based consensus protocol called Gasper (Buterin et al., 2020). Participants of Gasper are called validators. Each validator must deposit 32 ETH as collateral (Ethereum Foundation, 2024a). Time is divided into 12-second intervals called slots. This results in 7,200 slots in Ethereum per day. During each slot, a single validator, referred to as the proposer, proposes a block, while other validators vote for the proposed block if it extends the chain head in their local view.
@@
Proposer-builder separation (PBS) and MEV-Boost.   Proposer-builder separation (PBS) can be conceptually viewed as a sidecar to the core consensus protocol (although some proposals are more tightly coupled with consensus) to allow the proposer to outsource the job of building blocks to builders.
@@
MEV-Boost (Flashbots, 2024) is the current incarnation of the PBS idea. About 90% of all blocks on Ethereum are built by MEV-Boost (Wahrstätter, 2024). The basic idea is to auction off the right to propose a block to the builder who bids the highest in an open-bid ascending price auction (contributors, 2024). Because builders and proposers are mutually untrusted, the auction is facilitated by a (trusted) third party called a relay (Flashbots, 2022a). Here are the high-level steps (we refer readers to (Flashbots, 2024) for details):
@@
  • Builders submit bids to relays throughout the slot. For the purpose of this paper, we denote a bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key. Bids are signed by the builder but we omit the signature for clarity. Relays verify that $B$ indeed pays the proposer at least $BV$.
@@
  • The proposer selects a block hash $h_{B}$, typically the one with the highest $BV$, and sends the relay a signature on $h_{B}$. This signature is a promise to propose $B$ in the consensus protocol (if it proposes another block, this signature is evidence of equivocation and will cause the proposer to be penalized (Buterin et al., 2020).) After verifying the signature, the relay returns $B$ to the proposer.
Two additional remarks. First, relays are trusted entities, but anyone can run a relay (currently there are ten (EthStaker Community, 2024)), and builders and proposers can connect to the relays they trust. Second, winning bids are recorded on-chain, and losing bids are discarded after the auction. We collected and archived the historical bids over time to enable this study, as we will detail in Sec. 3.
2.2. MEV Supply Chain
@@
Wallet.   Users use wallets to interact with blockchains. A user communicates her intent (e.g., transferring cryptocurrency) to the wallet, which generates the transaction to fulfill the user’s intent and sends it to the network. MEV may be generated during this process, e.g., when the user transaction creates an arbitrage opportunity. While the wallet is typically controlled by the users, it can also be managed by a third party, such as Telegram trading bots (Intelligence, 2024). These bots are automated programs integrated with the Telegram app (Telegram, 2024), automatically conducting cryptocurrency trades on behalf of users. Users’ transactions eventually reach builders or proposers, either through the peer-to-peer network (the mempool) or through third-party services called private channels.
@@
Channel.   A private channel, or channel for short, connects users directly with builders, bypassing the mempool. Channels receive transactions from users and then, forward them to the builders specified by the users, offering privacy and atomicity guarantees. E.g., a user can send a bundle (list) of transactions and request them not to be unbundled. Examples include MEV-Share (Flashbots, 2024), MEV Blocker (MEV Blocker, 2023), BloXroute (bloXroute Labs, 2024), etc.
Searcher.   Searchers are a special type of users. They run (typically proprietary) algorithms to construct profitable transaction sequences (bundles) based on the pending transactions and the blockchain state. Same as regular users, searchers may submit transactions to mempool or channels. In addition, most builders have public APIs that allow searchers to submit bundles directly. Some searchers may prefer channels for the privacy and atomicity benefits. A search is said to integrate with a builder if it sends almost all of its orders to that builder exclusively. For example, searcher 0xA69b… is considered an integrated searcher of Beaver Builder because Beaver Builder includes the majority of its transactions.
@@
Private order flow.   Order flow refers to a source of transactions that may carry extractable value and can be categorized into two categories: public order flow, originating from the public mempool, and private order flow, sent directly by users or searchers without being broadcast to the public mempool. The entities that can provide private order flows to the builders are referred to as private order flow providers, which includes searchers, channels, users, and Telegram trading bots as shown in Fig. 2.
2.3. Model and Assumptions.
Accounts and states.   We follow the model of (Babel et al., 2023a, b). We use $\mathcal{A}$ to denote the space of all possible accounts. For example, in Ethereum, both user-owned and contract-owned accounts are represented by 160-bit identifiers. For $a\in\mathcal{A}$, $\text{balance}(a)$ denotes the balance of all tokens held in account $a$, and $\text{balance}(a)[\mathbf{T}]$ specifies the balance of token $\mathbf{T}$. For simplicity, $\text{balance}(a)[0]$ is used to denote the balance of the primary token (e.g., ETH in Ethereum).
@@
(1) $EV_{p}(B)=\sum_{a\in A_{p}}\text{balance}_{s^{\prime}}(a)[0]-\text{balance}_{s}(a)[0]$
@@
Builder’s income.   We assume a builder $p$’s income from building a block $B$ is $EV_{p}(B)$, or $EV(B)$ for short since the builder identity is usually encoded in the block. We compute EV following a similar approach as that in MEV analysis tools (EigenPhi, 2023; Flashbots, 2023b) and previous works (Heimbach et al., 2023; Thiery, 2023). Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees (Ethereum, 2024) and direct transfers (Flashbots, 2024a) to the builder’s addresses minus the refund to other addresses. We also refer to the builder’s income from a block (or a set of transactions in a block) as the value of the block (or the set of transactions).
@@
3. Dataset
We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to March 2024, and a dataset of full bids from April to August 2023. The full bids are obtained from the ultra sound relay (Money, 2023). Recall from Sec. 2.1 that a bid consists of a block $B$ and a bid value $BV$ and a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$ where $metadata$ includes parent hash, gas limit, gas used, etc.
To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below. Tab. 2 summarizes the datasets.
@@ 3. Dataset
On-chain blocks 4,019,895 Reth node
transactions 606,484,921
Auction blocks produced by MEV-Boost 3,517,851 relay API
partial bids 5,577,318,196 relay API
full bids from ultra sound 164,700,566 ultra sound relay
Misc. builder public keys 191 see Sec. 3.1
builder addresses 125
searcher addresses 2,069
private transactions 49,081,209
transaction sources 33,010,466
Table 2. Overview of collected datasets.
3.1. Data Collection and Validation
On-chain data.   We run an Ethereum execution client Reth (paradigmxyz, 2024) along with a consensus client Lighthouse (Prime, 2024) to collect blocks and transactions from September 15, 2022, to March 31, 2024. The on-chain dataset includes 4,019,895 blocks and 606,484,921 transactions.
Data from relay APIs.   We connected to all 13 known relays333three relays were inactive during the period of our study and collected the following information from their public APIs: 1) the winning bid including the final block sent to the proposer; and 2) partial bids for blocks that did not win the auction. Recall that partial bids do not include the block body. We collected 3,517,851 winning bids and 5,577,318,196 partial bids from September 15, 2022, to March 31, 2024, in a total of 3,810,630 auctions. We use winning bids to identify blocks produced through MEV-Boost. We use partial bids to quantify the importance of private order flow providers in Sec. 4.
To validate the completeness of the data collected from relay APIs, we compared the winning bids we collected with three public datasets (Network, 2023; Wahrstätter, 2023; DataAlways, 2024). Our dataset includes 5,382 more winning bids than other datasets combined. Upon comparison with on-chain blocks, we find they are all included on Ethereum, thereby indicating a high level of completeness.
Full bids from the ultra sound relay .   From the ultra sound relay (Money, 2023), we obtained the full bids in 147,926 historical auctions that took place between April and August 2023. Since the data volume is large (we record roughly 200 GB data per day), our dataset samples one week of each month: April 9-15, May 1-7, June 1-7, July 1-7, and August 1-7, 2023. The dataset consists of 164,700,566 full bids (block, bid value, and builder public key). We use full bids in Sec. 5 to quantify the competitiveness and efficacy of MEV-Boost auctions.
The original data from relays includes bids submitted after the deadline because of network latency. We removed them as we are only interested in the bids before the auction ends. On April 18, 2023, MEV-Boost introduced bid cancellation that allows a bidder to cancel a previous bid by submitting a lower one. However, proposers can query bids before they are canceled and store them locally (Ethereum Research, 2023), effectively bypassing cancellation. Therefore we do not take cancellation into consideration.
@@ 3.1. Data Collection and Validation
  • Builder public keys: We collect public keys controlled by each builder combining two data sources. First, some builders publish their public keys in official documents or code repositories. For example, Titan builder has disclosed 12 public keys in its official document (Builder, 2024). Second, some builders leave distinct marks in their blocks’ “extra field”. For example, Flashbots builder uses the mark ‘‘Illuminate Democratize Distribute’’. Given that such marks are unauthenticated and susceptible to impersonation, we ignore them unless they are used more than 100 times in the bids by the same builder public key.
@@
    To validate the completeness of our dataset, we manually cross-check it with public datasets (Rated Network, 2023; Network, 2024) and related works (Heimbach et al., 2024, 2023), and the results show that our dataset includes 62 more public keys, so our dataset is the most complete one so far.
@@
  • Builder addresses: We collect the builder addresses from the blocks’ last transactions combined with the public dataset (Etherscan, 2024). Similarly, to validate their completeness, we manually cross-check them with related works (Heimbach et al., 2024, 2023), and the results show that our dataset includes 113 more builder addresses, which is the most complete so far.
@@
  • Searcher addresses: We obtained known searcher addresses from Etherscan (Etherscan, 2024), libMEV (libMEV, 2024), and related work (Heimbach et al., 2024).
@@
  • Private transaction hashes: We use data from Mempool Guru (Yang et al., 2022) to label transactions that bypassed the public mempool as private. We believe these labels have already been cross-validated since the private transactions are collected by seven distributed nodes.
@@
  • Transaction sources: we are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through.
    We identified transactions originated from Telegram trading bots, including Maestro (Maestro, 2023a), Banana Gun (Banana Gun, 2023), and Unibot (Unibot X, 2023), by searching for transactions that interacted with their official smart contracts.
    Two channels, MEV Blocker and MEV-Share publish data (Protocol, 2024; Flashbots, 2024b) about transactions submitted to them. Specifically, transactions submitted to MEV Blocker are publicly available on Dune (Protocol, 2024), from which we extracted 14,901,303 transaction hashes that fall within our study’s range. We obtained 13,992,371 transactions submitted to MEV-Share from the public Flashbots dataset (Flashbots, 2024b). These transaction labels are sourced from their official datasets and we view them as ground truth.
4. Openess of the builder market
The builder market is technically open to everyone, but the barrier to accessing profitable private order flows (Sec. 2.2) forms an implicit access barrier.
To illustrate the importance of private order flow to the success of builders, Fig. 3 plots the fraction of builder income from private order flows. Currently, approximately 60% of the block value comes from private order flows.
Figure 3. Fraction of builder income from private order flows.
Finding I: The private order flow is crucial for builders to win the MEV-Boost auction because it contributes to more than 60% of the block value in over 50% of blocks.
Where can new builders get private order flows? The entities who provide (or sell) private order flows to builders are called private order flow providers. Different providers impose different access barriers to their private order flows. Some providers are open to new builders who satisfy certain requirements (such as MEV Blocker or MEV-Share), while some (e.g., most searchers) are almost completely opaque to the public.
To understand if the builder market is open to new builders, we first identify important “pivotal” providers and then quantify the cost for a new builder to access private order flows from representative pivotal providers.
4.1. Pivotal Private Order Flow Providers
We now present our method to quantify the importance of individual providers. An existing approach (Lu, 2024; Lu, Angela and Sui, Danning and Durnford, Jaden, 2024) is to compute the contribution of each provider to the builders’ income (Sec. 2.3) during a specific period. However, this approach only considers the magnitude of impact but neglects the duration (in terms of time or the number of auctions). For example, a single private transaction, such as one444 0xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip, could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot. Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that can affect half of the MEV-Boost auctions in over two weeks. This captures their sustained impact on the system.
For each pivotal provider, we compute the fraction of auctions in which it is pivotal. We call this metric the pivotal level, a number from 0 to 1. Fig. 4 plots the pivotal level of the top-5 providers.
Figure 4. A provider $P$ is pivotal if the winner of an auction could have lost without order flows from $P$. The pivotal level of top-5 pivotal providers over time. A line may not start from the beginning because the provider did not exist at that point. Maestro became an independent provider around December 2023.
Third, we observe a notable downtrend in the pivotal level of MEV-Share and MEV Blocker starting from November 2023, accompanied by an uptrend in Maestro, a popular Telegram trading bot (Maestro, 2023a), beginning in December 2023. This is because Maestro used to be an important upstream of MEV-Share and MEV Blocker but stopped in November 2023, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
Although platforms like MEV-Share and MEV Blocker positively affect decentralization, the incentives of their upstream providers may not be aligned. For example, Mastro may prefer to send flows to builders directly for a better price than what MEV-Share and MEV Blocker can offer. It is alarming to see that upstream providers can significantly swing the significance of platforms.
Finding II: We identified five pivotal providers (MEV-Share, MEV Blocker, jaredfromsubway.eth, Banana Gun, Maestro) who had a sustained influence on the outcome of more than half of MEV-Boost auctions. A positive trend towards decentralization is observed as the pivotal level of the searcher jaredfromsubway.eth decreased, while the pivotal level of MEV-Share and MEV Blocker increased. However, their reliance on upstream providers like Maestro is alarming.
Builder-specific Pivotal Level.    We compute each provider’s builder-specific pivotal level to understand the relationship between providers and builders. For a given provider $P$, its pivotal level specific to builder $B$ is the pivotal level of the MEV-Boost auctions where $B$ is the winner.
Fig. 5 plots selected providers’ builder-specific pivotal levels for the top-3 builders, ranked based on their weekly market share in April 2024. We include providers whose weekly pivotal levels exceed 25% for any of the three builders over two months to illustrate these builder-specific pivotal levels. We note a clear variation in the importance of providers for different builders. For example, the significance of MEV Blocker and Banana Gun to Titan Builder is greater than they are to other builders. Furthermore, we observe that some providers have a pivotal level exceeding 50% only with specific builders, suggesting a possible collaborative relationship. The searcher 0xA69b… consistently influences over 50% of the MEV-Boost auctions won by Beaver Builder, while the searchers 0x0087…, 0x2800…, and 0x51c7… successively affect those won by Rsync Builder. These pivotal providers are all integrated searchers of the corresponding builders that have been identified in previous works (Heimbach et al., 2024; Gupta et al., 2023). This observation corroborates findings from previous studies (Heimbach et al., 2024; Gupta et al., 2023; Thiery, 2023), confirming that the order flows from integrated searchers are crucial to those builders.
Figure 5. Weekly builder-specific pivotal level for top-3 builders.
4.2. Quantifying the Access Barrier
Pivotal providers have different degrees of openness to builders: searchers are generally not accessible, and platforms like MEV Share and MEV Blocker are more open, although they also set up certain barriers. In this section, we present more details about the access barrier set up by providers.
Searchers.   Since searchers are operated by anonymous entities without any public communication channels, there is no public statement about which builders may access their flows. Existing studies indicate that searchers tend to send their order flow to the top builders (Titan, 2023) or share it exclusively with vertically integrated builders (Titan, 2023; Thiery, 2023; Gupta et al., 2023).
Telegram trading bots.   The requirements for Telegram trading bots, in terms of sending order flow to builders, depend on whether they utilize channels like MEV Blocker (Maestro, 2023b; Labs, 2023). If so, their requirements are aligned with those of these channels. Banana Gun does not utilize these channels, nor does it explicitly state the requirements for builders.
The channels impose various requirements on a builder’s reputation. Market share—the percentage of blocks built by a given builder—is a metric most relay providers use to evaluate builders’ reputations. Intuitively, a builder with a high market share, seen as reputable, is less likely to compromise private order flows as doing so will forgo their future revenue. MeowRPC (Meow RPC, 2024) sends the order flow to trusted builders with high market share. Similarly, MEV-Share (Flashbots, 2024) requires that a builder possesses a competitive market share. MEV Blocker (MEV Blocker, 2023) requires that a builder must maintain at least 1% market share over one week.
The cost of establishing minimal market share.   New builders face a “chicken and egg” problem as they need access to private order flows to win auctions and gain market share, but the private order flow providers only serve builders with a decent market share. In practice, builders pay out of pocket to overbid in MEV Boost auctions, a practice known as block subsidization. This represents a capital entry barrier for new builders, which we are able to quantify using the auction dataset.
We compute the minimal subsidy needed to establish a 1% market share for one week (which is the requirement of MEV Blocker (MEV Blocker, 2023)) by replaying historical auctions with an added new builder who can only access the public order flow. We relegate the details to Sec. C.4.
Figure 6. Subsidy required to establish a 1% market share, as required by MEV Blocker. This quantifies the cost of accessing private order flows.
The result is plotted in Fig. 6. Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 1.4 ETH by March 2024. It is important to recognize that the costs derived from this method only represent the lower bound. Predicting the minimum cost of winning a MEV-Boost auction is challenging, so builders need to pay more to ensure victory. Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share. In reality, the subsidies new builders need to provide are substantially higher. For example, tbuilder (Tbuilder, 2024), a new builder in the builder market, subsidized 6.8 ETH in total for 645 blocks in the week before March 21, 2024 (Flashbots, 2024).
Finding III: The private order flow providers impose reputation requirements on the builders, which are usually evaluated based on market share. The minimum entry cost for new builders required to establish a required market share increases over time.
4.3. Trust Crisis and Centralization
Our findings highlight a trust issue in today’s MEV supply chain: the lack of trustworthy fair exchange mechanisms between parties creates order flow silos and harms decentralization.
Private order flow providers are at a disadvantage when interacting with builders and must trust builders not to unbundle the transactions. As a result, providers rely on an informal reputation system for self-protection, but it is not hard to see that reputation mechanisms are easily bypassed, as blockchain is fundamentally anonymous. Once a reputable builder receives the bundle from the searcher, the builder can create a Sybil and cheat the searcher (e.g., by unbundling the transaction bundle) under the new name.
While such an attack has not been reported in practice, it is noteworthy that builders occasionally unbundle searcher’s bundles. For example, the f1b builder unbundled a bundle from the searcher jaredfromsubway.eth at block height 18476515. Although the searcher was compensated afterward, this incident underscores the fragile security foundation of the trust chain in the MEV-Boost auction. We also noticed that one anonymous builder666public key: 0xad5ce1019364de39…35af575554a7aa6f might be the Sybil of a reputable builder because this builder’s blocks included order flows from sources that are only accessible by reputable builders.
The malicious actors within the trust chain pose a dual threat. On the one hand, security concerns deter searchers from submitting their bundles to multiple builders, further entrenching private order flow as a barrier to entry for new builders. On the other hand, integrating the roles of searcher and builder offers a potential mitigation to the security issue. Integrated searcher-builder has other structural advantages (Gupta et al., 2023; Pai and Resnick, 2023) (such as low latency) that may be unfair to regular builders.
In Sec. 7.2, we discuss directions to bridge the trust gap between searchers and builders.
Finding IV: The requirements for builders stem from the barrier of trust, as illustrated by the hidden trust crisis confirmed through real-world examples. Lowering or removing trust requirements in the MEV-Boost auction can facilitate decentralization.
5. Competition and Efficiency in the MEV-Boost auctions
Sec. 4 focuses on the evolution of builder market. In this section, we focus on the status quo and study the implications of a centralized builder market as we have today. First, we investigate the (in)equality of builders’ block-building capability, measured by the true values of builders’ bids in MEV-Boost auctions. Then, we quantify the competitiveness and efficiency of historical MEV Boost auctions. These two studies complement each other. The former evaluates whether today’s builder market has the potential to be competitive, while the latter evaluates whether the current mechanism in fact induces the desired outcomes. Our study includes some MEV-Boost auctions, the highest bids of which are not included on Ethereum due to network latency. However, we do not omit them since the auctions themselves are still meaningful.
5.1. Inequality in Block-building Capacity
A primary metric to evaluate a builder’s block-building capacity is the MEV it extracts over time. In MEV Boost auctions, this quantity is captured by the true values of the builder’s bids.

### v3

@@ 2.1 MEV-Boost Auctions
MEV.   Miner/Maximal Extractable Value (MEV) [16] refers to the profit that privileged players (e.g., validators) can earn by including, excluding, and reordering the transactions. MEV poses a threat to decentralization, as explained in the introduction. There are two schools of thought on mitigating the negative impacts of MEV. The first one is referred to as MEV democratization [30], the main idea of which is to facilitate MEV extraction and level the playground. MEV-Boost [35], to be introduced below, is the primary example. Another approach is to enforce certain ordering policies (e.g., FIFO) to prevent order manipulation, such as [45, 44]. We focus on MEV-Boost because it is widely used.
@@
PoS Ethereum.   Before introducing background on MEV-Boost auctions, we first cover the relevant concepts in Ethereum’s Proof of Stake-based consensus protocol called Gasper [13]. Participants of Gasper are called validators. Each validator must deposit 32 ETH as collateral [22]. Time is divided into 12-second intervals called slots. This results in 7,200 slots in Ethereum per day. During each slot, a single validator, referred to as the proposer, proposes a block, while other validators vote for the chain head in their local view.
@@
Proposer-builder separation (PBS) and MEV-Boost.   PBS can be conceptually viewed as a sidecar to the core consensus protocol (although some proposals are more tightly coupled with consensus) to allow the proposer to outsource the job of building blocks to builders.
@@
MEV-Boost [35] is the current incarnation of the PBS idea. About 90% of all blocks on Ethereum are built by MEV-Boost [87]. The basic idea is to auction off the right to propose a block to the builder who bids the highest in an open-bid ascending price auction [14]. Because builders and proposers are mutually untrusted, the auction is facilitated by a (trusted) third party called a relay [29]. Here are the high-level steps (we refer readers to [35] for details):
@@
  • Builders submit bids to relays throughout the slot. For this paper, we denote a bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key. Bids are signed by the builder but we omit the signature for clarity. Relays verify that $B$ indeed pays the proposer at least $BV$.
@@
  • The proposer selects a block hash $h_{B}$, typically the one with the highest $BV$, and sends the relay a signature on $h_{B}$. This signature is a promise to propose $B$ in the consensus protocol (if it proposes another block, this signature is evidence of equivocation and will cause the proposer to be penalized [13].) After verifying the signature, the relay returns $B$ to the proposer.
Two additional remarks. First, relays are trusted entities, but anyone can run a relay (currently ten [27]), and builders and proposers can connect to the relays they trust. Second, winning bids are recorded on-chain, and losing bids are discarded after the auction. We collected and archived the historical bids over time to enable this study, as we will detail in Sec. 3.
2.2 MEV Supply Chain
@@
Wallet, users, trading bots.   Users use wallets to interact with blockchains. A user communicates her intent (e.g., transferring cryptocurrency) to the wallet, which generates the transaction to fulfill the user’s intent and sends it to the network. MEV may be generated during this process, e.g., when the user transaction creates an arbitrage opportunity. While the wallet is typically controlled by the users, it can also be managed by a third party, such as Telegram trading bots [43]. These bots are automated programs integrated with the Telegram app [79], automatically conducting cryptocurrency trades on behalf of users. Users’ transactions eventually reach builders or proposers, either through the peer-to-peer network (the mempool) or through third-party services called private channels.
@@
Channel.   A private channel, or channel for short, connects users directly with builders, bypassing the mempool. Channels receive transactions from users and then, forward them to the user-specified builders, offering privacy and atomicity guarantees. E.g., a user can send a bundle (list) of transactions and request them not to be unbundled. Examples include MEV-Share [37], MEV Blocker [60], BloXroute [9], etc.
Searcher.   Searchers are a special type of users. They run (typically proprietary) algorithms to construct profitable transaction sequences (bundles) based on the pending transactions and the blockchain state. Same as regular users, searchers may submit transactions to mempool or channels. In addition, most builders have public APIs that allow searchers to submit bundles directly. Some searchers may prefer channels for the privacy and atomicity benefits. A searcher is said to integrate with a builder if it sends almost all of its orders to that builder exclusively. For example, searcher 0xA69b… is considered an integrated searcher of Beaver Builder because Beaver Builder includes the majority of its transactions.
@@
Private order flow.   Order flow refers to a stream of transactions that may carry extractable value and can be categorized into two categories: public order flow from the public mempool, and private order flow, sent directly by users or searchers without being broadcast to the mempool. The entities that can provide private order flows to the builders are referred to as private order flow providers, which include searchers, channels, users, and Telegram trading bots as shown in Fig. 3.
2.3 Model and Assumptions.
Accounts and states.   We follow the model of [2, 3]. We use $\mathcal{A}$ to denote the space of all possible accounts. For example, in Ethereum, both user-owned and contract-owned accounts are represented by 160-bit identifiers. For $a\in\mathcal{A}$, $\text{balance}(a)$ denotes the balance of all tokens held in account $a$, and $\text{balance}(a)[\mathbf{T}]$ specifies the balance of token $\mathbf{T}$. For simplicity, $\text{balance}(a)[0]$ is used to denote the balance of the primary token (e.g., ETH in Ethereum).
@@
$EV_{p}(B)=\sum_{a\in A_{p}}\text{balance}_{s^{\prime}}(a)[0]-\text{balance}_{s}(a)[0]$ (1)
@@
Builder’s income.   We assume a builder $p$’s income from building a block $B$ is $EV_{p}(B)$, or $EV(B)$ for short since the builder identity is usually encoded in the block. We compute EV following a similar approach as that in MEV analysis tools [20, 31] and previous works [41, 80]. Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees [21] and direct transfers [32] to the builder’s addresses minus the refund to other addresses. We also refer to the builder’s income from a block (or a set of transactions in a block) as the value of the block (or the set of transactions).
@@
3 Dataset
We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to July 2024, and a dataset of full bids from April to December 2023. The full bids are obtained from the ultra sound relay [62]. Recall from Sec. 2.1 that a bid consists of a block $B$ and a bid value $BV$ and a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$ where $metadata$ includes parent hash, gas limit, gas used, etc. To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below. Tab. 1 summarizes the datasets.
@@ 3 Dataset
On-chain blocks 4,892,579 Reth node
transactions 748,859,795
Auction blocks produced by MEV-Boost 4,305,571 relay API
partial bids 8,789,557,915 relay API
full bids from ultra sound 326,574,676 ultra sound relay
Misc. builder public keys 248 see Sec. 3.1
builder addresses 102
searcher addresses 2,140
private transactions 77,747,623
transaction sources 60,345,066
Table 1: Overview of collected datasets.
3.1 Data Collection and Validation
On-chain data.   We run an Ethereum execution client Reth [68] along with a consensus client Lighthouse [69] to collect blocks and transactions from September 15, 2022, to July 31, 2024. The on-chain dataset includes 4,892,579 blocks and 748,859,795 transactions.
Data from relay APIs.   We connected to all 13 known relays [^2] Three relays were inactive during the period of our study. and collected the following information from their public APIs: 1) the winning bid including the final block sent to the proposer; and 2) partial bids for blocks that did not win the auction. Recall that partial bids do not include the block body. We collected 4,305,571 winning bids and 8,789,557,915 partial bids from September 15, 2022, to July 31, 2024. We use winning bids to identify blocks produced through MEV-Boost. We use partial bids to quantify the importance of private order flow providers in Sec. 5.
To validate the completeness of the data collected from relay APIs, we compared the winning bids we collected with three public datasets [64, 85, 17]. Our dataset includes 720 more winning bids than other datasets combined. Upon comparison with on-chain blocks, we find they are all included on Ethereum, thereby indicating a high level of completeness.
Full bids from the ultra sound relay .   From the ultra sound relay [62], we obtained the full bids in 301,479 historical auctions that took place between April and December 2023 (ultra sound relay stopped collecting bids at the end of 2023). Since the data volume is large (we record roughly 200 GB of data per day), our dataset samples one week from each month: April 9-15 and May 1-7 to December 1-7, 2023. The dataset comprises 326,574,676 full bids (block, bid value, and builder public key). We use full bids in Sec. 4 to quantify the competitiveness and efficiency of MEV-Boost auctions.
The original data from relays includes bids submitted after the deadline because of network latency. We removed them as we are only interested in the bids before the auction ends. On April 18, 2023, MEV-Boost introduced bid cancellation that allows a bidder to cancel a previous bid by submitting a lower one. However, proposers can query bids before they are canceled and store them locally [24], effectively bypassing cancellation. Therefore we do not consider cancellation.
@@ 3.1 Data Collection and Validation
  • Builder public keys: We collect public keys controlled by each builder combining two data sources: the public keys in builders’ official documents; and the distinct marks in their blocks’ “extra field” (see Appx. B for details).
@@
    To validate the completeness of our dataset, we manually cross-check it with public datasets [72, 65] and related works [42, 41], and the results show that our dataset includes 119 more public keys, so our dataset is the most complete one so far.
@@
  • Builder addresses: We collect the builder addresses from the blocks’ last transactions combined with the public dataset [26]. Similarly, to validate their completeness, we manually cross-check them with related works [42, 41], and the results show that our dataset includes 86 more builder addresses, which is the most complete so far.
@@
  • Searcher addresses: We obtained known searcher addresses from Etherscan [25], libMEV [50], and related work [42].
@@
  • Private transaction hashes: We use data from Mempool Guru [92] to label transactions that bypassed the public mempool as private. We believe these labels have already been cross-validated since the private transactions are collected by seven distributed nodes.
@@
  • Transaction sources: we are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through. We identified transaction sources using ground truth data (as described in Appx. B).
4 Implications of Builder Centralization
In this section, we focus on the status quo and study the implications of a centralized builder market as we have today. In particular, we are interested in the losses incurred by proposers, as significant losses could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
We quantify proposers’ losses in historical MEV-Boost auctions considering two factors: the competitiveness of the auctions and the inequality in block-building capacity.
Our study includes some MEV-Boost auctions where the highest bids were not included on Ethereum due to network latency. However, we do not omit these auctions since the auctions themselves remain meaningful.
4.1 Competitiveness of MEV-Boost Auctions
Ideally, revenue from MEV-Boost auctions should be near optimal due to competition. However, whether real-world auctions induce such near-perfect competition is complicated by several confounding factors: builders may not be able to respond to bids due to network latency, or they may collude implicitly or explicitly. In this section, we quantify the competitiveness of the MEV-Boost auctions using what we call competitive index (CI).
We also look at the allocation efficiency as it evaluates whether the current mechanism induces the desired outcomes.